# Accelerating Urban Science Research with AI Urban Scientist

### v1

@@ Abstract
Figure 6: Workflow of the Automatic Data Analysis Agent. The agent receives a hypothesis–dataset pair and generates an experiment plan by retrieving relevant examples from a code base of 10K+ scripts. It then produces executable multi-language code, runs the scripts in controlled environments, and iteratively debugs errors. When empirical data are insufficient, the agent can call external simulators to generate synthetic evidence. Together, these capabilities enable end-to-end empirical testing for heterogeneous urban science tasks.
Urban science requires the integration of highly heterogeneous datasets and the execution of diverse analytical tasks, ranging from high-dimensional statistical modelling to machine learning, spatial analysis, simulation, and causal inference. These tasks often involve multiple programming languages (e.g., Python, R) and specialised software environments. However, existing data–science agents are far from capable of handling such complexity: they can neither design multi-step analysis pipelines nor reliably generate executable code across heterogeneous tools and data modalities. To address this gap, we introduce an Automatic Data Analysis Agent (Figure 6) that performs end-to-end empirical analysis for a given hypothesis–dataset pair.
@@
Rather than relying solely on prompting or ad-hoc generation, the agent is grounded in a code knowledge base containing more than 10,000 scripts extracted from published urban science papers. These scripts cover common analytical tasks such as regression models, spatiotemporal analysis, exposure–response modelling, machine learning prediction, uncertainty quantification, and robustness checks. By retrieving relevant examples from this code base through RAG (retrieval-augmented generation), the agent constructs an experiment plan that outlines the analytical steps required to test the hypothesis.
@@
In many urban science problems, empirical data may be missing, insufficient, or unsuitable for directly testing a hypothesis. To address such cases, the agent can call external simulators—such as climate–health models, mobility simulators, agent-based city models, or environmental exposure models—by generating the required configuration files and code wrappers. The simulator output is then integrated into the analysis pipeline, enabling counterfactual reasoning or synthetic experiments when observational data are limited.
@@
Through these stages—knowledge-driven planning, retrieval-based code guidance, multi-language code generation, error correction, and simulator integration—the Automatic Data Analysis Agent provides a practical and robust mechanism for end-to-end empirical urban science research.
@@ Discussion
Lack of systematic evaluation benchmarks

### v2

@@ Abstract
Figure 4: Workflow of the Automatic Data Analysis Agent. The agent receives a hypothesis–dataset pair and generates an experiment plan by retrieving relevant examples from a code base of 10K+ scripts. It then produces executable multi-language code, runs the scripts in controlled environments, and iteratively debugs errors. When empirical data are insufficient, the agent can call external simulators to generate synthetic evidence. Together, these capabilities enable end-to-end empirical testing for heterogeneous urban research tasks.
urban research requires the integration of highly heterogeneous datasets and the execution of diverse analytical tasks, ranging from high-dimensional statistical modelling to machine learning, spatial analysis, simulation, and causal inference. These tasks often involve multiple programming languages (e.g., Python, R) and specialised software environments. However, existing data–science agents are far from capable of handling such complexity: they can neither design multi-step analysis pipelines nor reliably generate executable code across heterogeneous tools and data modalities. To address this gap, we introduce an Automatic Data Analysis Agent (Figure 4) that performs end-to-end empirical analysis for a given hypothesis–dataset pair.
@@
Rather than relying solely on prompting or ad-hoc generation, the agent is grounded in a code knowledge base containing more than 10,000 scripts extracted from published urban research papers. These scripts cover common analytical tasks such as regression models, spatiotemporal analysis, exposure–response modelling, machine learning prediction, uncertainty quantification, and robustness checks. By retrieving relevant examples from this code base through RAG (retrieval-augmented generation), the agent constructs an experiment plan that outlines the analytical steps required to test the hypothesis.
@@
In many urban research problems, empirical data may be missing, insufficient, or unsuitable for directly testing a hypothesis. To address such cases, the agent can call external simulators—such as climate–health models, mobility simulators, agent-based city models, or environmental exposure models—by generating the required configuration files and code wrappers. The simulator output is then integrated into the analysis pipeline, enabling counterfactual reasoning or synthetic experiments when observational data are limited.
@@
Through these stages—knowledge-driven planning, retrieval-based code guidance, multi-language code generation, error correction, and simulator integration—the Automatic Data Analysis Agent provides a practical and robust mechanism for end-to-end empirical urban research research.
Reliable Urban Research Critic
To ensure that the hypotheses generated by the Ideation Agent are evaluated through a lens consistent with the standards of high-quality urban research, we develop a Critic Agent that emulates the judgment patterns of expert reviewers in the field. Unlike general-purpose AI reviewers, which tend to inherit evaluation criteria from machine learning venues such as OpenReview, our Critic Agent is explicitly grounded in the epistemic standards of Nature and Nature Cities, two flagship journals representing the highest level of urban and interdisciplinary research. As shown in Figure 2, the key steps of constructing the Critic Agent are as follows.
System-Level Alignment with urban research Standards.
We begin by constructing a domain-specific system prompt that integrates the editorial requirements, aims and scope, and mission statements of Nature and Nature Cities. These statements emphasize conceptual significance, interdisciplinary depth, empirical grounding, methodological rigor, and societal relevance. By embedding these principles directly into the system prompt, we ensure that the Critic Agent evaluates hypotheses according to norms that are aligned with the expectations of top-tier urban research journals.
Fine-Tuning with Expert Review Corpora.
Impact-Factor–Aligned Evaluation Calibration.
Through system prompt alignment, expert-review fine-tuning, and impact-factor–based calibration, the Critic Agent develops evaluation capabilities that reflect the values of the urban research community. It can judge whether a hypothesis has conceptual significance, cross-disciplinary relevance, empirical feasibility, and theoretical credibility. Its output includes both a tier rating (Tier 1–4 or reject) and explanatory feedback, enabling the hypothesis refinement loop to converge toward scientifically meaningful ideas.
Figure 5: The construction and training pipeline of the Critic Agent. We align the Critic Agent with domain standards by integrating the editorial principles of Nature and Nature Cities as the system prompt. A corpus of 15K papers provides 15K+ seed hypotheses, while 2K+ expert reviewer comments from Nature family journals supply high-quality positive and negative evaluation signals for fine-tuning. Additionally, hypotheses extracted from the 15K papers are labeled into four tiers according to journal impact factor, enabling multi-level calibration. Together, these components produce a domain-informed Critic Agent capable of assigning idea-quality tiers (Tier 1–4 or reject) consistent with expert judgment in urban research.
@@
Building systematic evaluation benchmarks